What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 9% of par, and a current market price of (a) $68, (b) $81, (c) $117, and (d) $134? Round your answers to two decimal places. a. b. C. d.
Answer #1

Annualdividend=$100*9%=$9

Nominal rate=Annual dividend/Current market price

Case Nominal rate
a (9/68)=13.24%(Approx)
b (9/81)=11.11%(Approx)
c (9/117)=7.69%(Approx)
d (9/134)=6.72%(Approx)